# ZipCache: A DRAM/SSD Cache with Built-in Transparent Compression

### v1

@@ 4 Evaluation
Table 2: Four categories of workload locality.
@@
To study the impact of cache object content compressibility, we modified Cachebench to generate the cache object content as follows. Given a parameter $\eta\in[0\%,100\%)$, we fill $1-\eta$ of each cache object with incompressible random content and set the remaining $\eta$ portion as all zeros. Hence, the cache object content compressibility improves as $\eta$ increases. The same method has been used by the popular I/O test tool FIO (flexible I/O tester) [3] to generate I/O data with configurable compressibility.
4.1 Overall Cache Performance
We first evaluate and compare the speed performance of CacheLib [2], Kangaroo [46] (a variant of CacheLib for reducing SSD write amplification), xcache [11], and ZipCache. To cover a wide range of spectrum, we considered the scenarios when the total active working set size is either larger or smaller than the DRAM/SSD cache capacity. For both scenarios, we set the key size as 16B and cache object size as 64B. The cache object content is generated with the compressibility parameter $\eta=50\%$. The workload consists of 16 user threads issuing GET and PUT requests at the ratio of 1:1. Since adaptive compression bypassing is effective only under workloads with very strong locality, we disable this feature here an will study its effect later in Section 4.4. The parameters of CacheLib and xcache are set as their default values, and we turned on the LZ4 compression of xcache’s SSD cache tier. For ZipCache, the leaf page size of DRAM tier B+ tree and SSD tier B+ tree is set to 4KB and 64KB, respectively, and the DRAM tier per-page write buffer size limit is set to 256B.
@@ 4.1 Overall Cache Performance
4.2 DRAM Cache Tier Compression
@@
Figure 9: GET latency under workloads with moderate locality and different $C_{DRAM}:C_{WS}$.
@@
Figure 10: Average throughput of PUT-only and GET-only workload with moderate locality and different $C_{DRAM}:C_{WS}$.
@@
4.3 SSD Cache Tier Write Amplification
We further compared the SSD cache tier write amplification among ZipCache, CacheLib [2], Kangaroo [46], and xcache [11]. As discussed above in Section 3.3, once we deploy a hybrid-DRAM/SSD cache on SSDs with built-in compression, we can express the overall SSD write amplification as $WA=WA_{host}/WR_{NAND}$, where the host-side write amplification $WA_{host}\geq 1$ represents the write amplification induced by host-side cache software and intra-SSD write reduction $WR_{NAND}\geq 1$ quantifies the effect of intra-SSD compression (i.e., compression ratio achieved by SSD).
@@ 4.3 SSD Cache Tier Write Amplification
Figure 11: (a) Host-side write amplification $WA_{host}$ comparison among CacheLib, Kangaroo and ZipCache under different cache object size, and (b) intra-SSD write reduction under different per-4KB fill-factor (xcache is not included since its intra-SSD write reduction remains as 1).
Fig. 11(a) shows the host-side write amplification $WA_{host}$ under the Cachebench workload with moderate locality. The key size is 16B and cache object size is 32B, 64B, and 128B. By directly hashing each cache object into one SSD 4KB LBA, CacheLib experiences very high $WA_{host}$ that is inversely proportional to the cache object size. For example, its $WA_{host}$ increases from 31.5 to 60.7 when the cache object size reduces from 128B to 64B. Regarding xcache, its DRAM cache uses hash index and SSD cache uses LSM-tree index. Hence, xcache has lower SSD write amplification than CacheLib. By complementing CacheLib with a WAL to accumulate multiple cache objects hashed to the same SSD 4KB LBA, Kangaroo [46] reduces $WA_{host}$ at the cost of SSD storage capacity. Let $\delta<1$ denote the ratio of WAL size and SSD cache size, we considered two values of $\delta$: 5% and 10%. Fig. 11(a) shows the effectiveness of WAL-assisted write amplification and the trade-off between the WAL-induced storage overhead $\delta$ and host-side write amplification $WA_{host}$. For ZipCache, we considered three leaf page size of SSD tier B+ tree $BT_{SSD}$, including 4KB, 16KB, and 64KB. As discussed above, to minimize SSD tier cache hit time, $BT_{SSD}$ keeps all its non-leaf pages in DRAM and leaves leaf pages on SSD. We define the $BT_{SSD}$ memory overhead (denoted as $\zeta$) as the ratio between the total size of its in-memory non-leaf pages and total size of its on-SSD leaf pages. The $BT_{SSD}$ leaf page size affects the trade-off between host-side write amplification $WA_{host}$ and $BT_{SSD}$ memory overhead $\zeta$, which can be observed from Fig. 11(a) and Table 3. The results show that, dependent upon their different configurations, Kangaroo and ZipCache have comparable host-side write amplification, which is slightly better than xcache and significantly better than that of CacheLib, especially under small cache object size.
Table 3: $BT_{SSD}$ memory overhead $\zeta$.
@@
4.4 Adaptive Compression Bypassing
@@
Figure 12: (a) GET latency and (b) average throughput comparison under Cachebench workload with strong locality, where GET:PUT ratio is 30%:70% or 70%:30%.
@@
Figure 13: Change of the average GET latency after the sudden hot region position shift.
4.5 Sensitivity Study
@@
Figure 14: (a) DRAM tier cache hit ratio and (b) overall throughput under different compressibility parameter $\eta$.
Figure 15: Throughput and memory usage overhead under different per-page write buffer size, where PUT-only Cachebench workload with strong locality is used.
@@
Figure 16: GET latency for various sizes objects under the Cachebench workload with modest locality.
5 Related Work
@@
Key-Value Stores: The design and implementation of in-memory and SSD-based key-value stores have been extensively studied by the research community. Most in-memory KV stores chose to employ hash-based index (e.g., Memcached [4], Redis [6], MemC3 [31], MICA [43], Mega-KV [61], RAMCloud [49], FASTER [20], HotRing [22]). Only few implementations, such as MassTree [44], use tree-based index. SSD-based key-value stores typically employ tree-based index and support block data compression. Because of its low write amplification, log-structured merge (LSM) tree [50] index data structure has received most recent attentions in research community for building SSD-based key-value stores [54, 53, 16, 26, 35, 21].
@@
Prior work also studied the design of hybrid key-value stores over different memory technologies (e.g., DRAM and SSD), aiming to strike a better balance between the speed and cost. CacheLib [18] and xcache [11] are two representative hybrid-DRAM/SSD key-value stores, where CacheLib uses hash-based index for both the DRAM and SSD tiers and xcache uses hash-based index for DRAM tier and LSM-tree index for SSD tier. Motivated by intensive research on NVM (non-volatile memory) technologies over the past decade, recent work has developed NVM-enhanced hybrid key-value stores [60, 59, 27, 29]. Though there is prior research on flash cache compression [42, 56, 41, 34], little work studied the potential of improving cache performance via in memory block data compression.
@@
Memory Compression: The computer architecture research community has widely studied the implementation of hardware-based main memory compression [24, 30, 51, 62]). Aiming at better serving general-purpose computing systems, hardware-based main memory compression focuses on fine-grained per-cacheline (e.g., 64B) compression. The Linux kernel feature Zswap [12] compresses to-be-swapped 4KB pages and keeps them in DRAM, which has been used by Google [38] and Meta [57] to increase effective DRAM capacity. Contemporary data analytics systems like SAP HANA [32], Oracle [39], and Snowflake [25] apply block compression to their in-memory column-stores to reduce their memory consumption. In-memory time series databases [13, 52] also widely use compression to exploit the inherently high compressibility of time series data.
@@
Computational SSD: Computational storage has attracted significant recent interest [19, 17, 37, 55], and commercial products are emerging on the commercial market (e.g., Samsung’s SmartSSD [7] and ScaleFlux’s CSD [8]). Recent research [63, 23, 36] has studied how database management systems could take advantage of computational SSDs with built-in transparent compression.
6 Conclusion

### v2

@@ 4. Evaluation
Table 2. Four categories of workload locality.
@@
To study the impact of cache object content compressibility, we modified Cachebench to generate the cache object content as follows. Given a parameter $\eta\in[0\%,100\%)$, we fill $1-\eta$ of each cache object with incompressible random content and set the remaining $\eta$ portion as all zeros. Hence, the cache object content compressibility improves as $\eta$ increases. The same method has been used by the popular I/O test tool FIO (flexible I/O tester) (FIO, 2023) to generate I/O data with configurable compressibility.
4.1. Overall Cache Performance
Figure 7. Comparison of throughput and overall DRAM/SSD cache hit ratio when the active working set is much larger than the DRAM/SSD cache capacity. We fixed the total active working set size $C_{WS}$ as 6TB and, under each category of workload locality, considered four different settings of {$C_{DRAM}$, $C_{SSD}$}: {64GB, 0.5TB}, {64GB, 1TB}, {128GB, 0.5TB}, and {128GB, 1TB}.
Figure 8. Comparison of throughput and DRAM cache hit ratio when the active working set fits into the DRAM/SSD cache (hence the SSD cache tier hit ratio is 100%). We fixed the DRAM cache tier capacity $C_{DRAM}$ as 64GB and considered $C_{DRAM}:C_{WS}$ ratio of 1:16, 1:8, 1:4, and 1:2.
We first evaluate and compare the speed performance of CacheLib (Cac, 2024), Kangaroo (McAllister et al., 2021) (a variant of CacheLib for reducing SSD write amplification), xcache (xca, 2022), and ZipCache. To cover a wide range of spectrum, we considered the scenarios when the total active working set size is either larger or smaller than the DRAM/SSD cache capacity. For both scenarios, we set the key size as 16B and cache object size as 64B. The cache object content is generated with the compressibility parameter $\eta=50\%$. The workload consists of 16 user threads issuing GET and PUT requests at the ratio of 1:1. Since adaptive compression bypassing is effective only under workloads with very strong locality, we disable this feature here an will study its effect later in Section 4.4. The parameters of CacheLib and xcache are set as their default values, and we turned on the LZ4 compression of xcache’s SSD cache tier. For ZipCache, the leaf page size of DRAM tier B+ tree and SSD tier B+ tree is set to 4KB and 64KB, respectively, and the DRAM tier per-page write buffer size limit is set to 256B.
@@ 4.1. Overall Cache Performance
4.2. DRAM Cache Tier Compression
@@
Figure 9. GET latency under workloads with moderate locality and different $C_{DRAM}:C_{WS}$.
@@
Figure 10. Average throughput of PUT-only and GET-only workload with moderate locality and different $C_{DRAM}:C_{WS}$.
@@
4.3. SSD Cache Tier Write Amplification
We further compared the SSD cache tier write amplification among ZipCache, CacheLib (Cac, 2024), Kangaroo (McAllister et al., 2021), and xcache (xca, 2022). As discussed above in Section 3.3, once we deploy a hybrid-DRAM/SSD cache on SSDs with built-in compression, we can express the overall SSD write amplification as $WA=WA_{host}/WR_{NAND}$, where the host-side write amplification $WA_{host}\geq 1$ represents the write amplification induced by host-side cache software and intra-SSD write reduction $WR_{NAND}\geq 1$ quantifies the effect of intra-SSD compression (i.e., compression ratio achieved by SSD).
@@ 4.3. SSD Cache Tier Write Amplification
Figure 11. (a) Host-side write amplification $WA_{host}$ comparison among CacheLib, Kangaroo and ZipCache under different cache object size, and (b) intra-SSD write reduction under different per-4KB fill-factor (xcache is not included since its intra-SSD write reduction remains as 1).
Fig. 11(a) shows the host-side write amplification $WA_{host}$ under the Cachebench workload with moderate locality. The key size is 16B and cache object size is 32B, 64B, and 128B. By directly hashing each cache object into one SSD 4KB LBA, CacheLib experiences very high $WA_{host}$ that is inversely proportional to the cache object size. For example, its $WA_{host}$ increases from 31.5 to 60.7 when the cache object size reduces from 128B to 64B. Regarding xcache, its DRAM cache uses hash index and SSD cache uses LSM-tree index. Hence, xcache has lower SSD write amplification than CacheLib. By complementing CacheLib with a WAL to accumulate multiple cache objects hashed to the same SSD 4KB LBA, Kangaroo (McAllister et al., 2021) reduces $WA_{host}$ at the cost of SSD storage capacity. Let $\delta<1$ denote the ratio of WAL size and SSD cache size, we considered two values of $\delta$: 5% and 10%. Fig. 11(a) shows the effectiveness of WAL-assisted write amplification and the trade-off between the WAL-induced storage overhead $\delta$ and host-side write amplification $WA_{host}$. For ZipCache, we considered three leaf page size of SSD tier B+ tree $BT_{SSD}$, including 4KB, 16KB, and 64KB. As discussed above, to minimize SSD tier cache hit time, $BT_{SSD}$ keeps all its non-leaf pages in DRAM and leaves leaf pages on SSD. We define the $BT_{SSD}$ memory overhead (denoted as $\zeta$) as the ratio between the total size of its in-memory non-leaf pages and total size of its on-SSD leaf pages. The $BT_{SSD}$ leaf page size affects the trade-off between host-side write amplification $WA_{host}$ and $BT_{SSD}$ memory overhead $\zeta$, which can be observed from Fig. 11(a) and Table 3. The results show that, dependent upon their different configurations, Kangaroo and ZipCache have comparable host-side write amplification, which is slightly better than xcache and significantly better than that of CacheLib, especially under small cache object size.
Table 3. $BT_{SSD}$ memory overhead $\zeta$.
@@
4.4. Adaptive Compression Bypassing
@@
Figure 12. (a) GET latency and (b) average throughput comparison under Cachebench workload with strong locality, where GET:PUT ratio is 30%:70% or 70%:30%.
@@
Figure 13. Change of the average GET latency after the sudden hot region position shift.
4.5. Sensitivity Study
@@
Figure 14. (a) DRAM tier cache hit ratio and (b) overall throughput under different compressibility parameter $\eta$.
Figure 15. Throughput and memory usage overhead under different per-page write buffer size, where PUT-only Cachebench workload with strong locality is used.
@@
Figure 16. GET latency for various sizes objects under the Cachebench workload with modest locality.
5. Related Work
@@
Key-Value Stores: The design and implementation of in-memory and SSD-based key-value stores have been extensively studied by the research community. Most in-memory KV stores chose to employ hash-based index (e.g., Memcached (Mem, 2018), Redis (Red, 2023), MemC3 (Fan et al., 2013), MICA (Lim et al., 2014), Mega-KV (Zhang et al., 2015), RAMCloud (Ousterhout et al., 2015), FASTER (Chandramouli et al., 2018), HotRing (Chen et al., 2020)). Only few implementations, such as MassTree (Mao et al., 2012), use tree-based index. SSD-based key-value stores typically employ tree-based index and support block data compression. Because of its low write amplification, log-structured merge (LSM) tree (O’Neil et al., 1996) index data structure has received most recent attentions in research community for building SSD-based key-value stores (Roc, 2023; Raju et al., 2017; Balmau et al., 2017; Dayan and Idreos, 2018; Huang et al., 2019; Chen et al., 2021a).
@@
Prior work also studied the design of hybrid key-value stores over different memory technologies (e.g., DRAM and SSD), aiming to strike a better balance between the speed and cost. CacheLib (Berg et al., 2020) and xcache (xca, 2022) are two representative hybrid-DRAM/SSD key-value stores, where CacheLib uses hash-based index for both the DRAM and SSD tiers and xcache uses hash-based index for DRAM tier and LSM-tree index for SSD tier. Motivated by intensive research on NVM (non-volatile memory) technologies over the past decade, recent work has developed NVM-enhanced hybrid key-value stores (Zhan et al., 2020; Yao et al., 2020; Ding et al., 2022; Duan et al., 2023). Though there is prior research on flash cache compression (Li et al., 2016; Wang et al., [n. d.]; Li et al., [n. d.]; Hao et al., [n. d.]), little work studied the potential of improving cache performance via in memory block data compression.
@@
Memory Compression: The computer architecture research community has widely studied the implementation of hardware-based main memory compression (Choukse et al., 2018; Ekman and Stenstrom, 2005; Pekhimenko et al., 2013; Zhao et al., 2015)). Aiming at better serving general-purpose computing systems, hardware-based main memory compression focuses on fine-grained per-cacheline (e.g., 64B) compression. The Linux kernel feature Zswap (Zsw, 2024) compresses to-be-swapped 4KB pages and keeps them in DRAM, which has been used by Google (Lagar-Cavilla et al., 2019) and Meta (Weiner et al., 2022) to increase effective DRAM capacity. Contemporary data analytics systems like SAP HANA (Färber et al., 2012), Oracle (Lahiri et al., 2015), and Snowflake (Dageville et al., 2016) apply block compression to their in-memory column-stores to reduce their memory consumption. In-memory time series databases (Adams et al., 2020; Pelkonen et al., 2015) also widely use compression to exploit the inherently high compressibility of time series data.
@@
Computational SSD: Computational storage has attracted significant recent interest (Cao et al., 2020; Barbalace and Do, 2021; Kwon et al., 2021; Vinçon et al., 2022), and commercial products are emerging on the commercial market (e.g., Samsung’s SmartSSD (Sma, 2024) and ScaleFlux’s CSD (Sca, 2024)). Recent research (Zheng et al., 2020; Chen et al., 2021b; Huang et al., 2023) has studied how database management systems could take advantage of computational SSDs with built-in transparent compression.
6. Conclusion
@@
Acknowledgements.
This work was supported in part by the National Science Foundation (NSF) under grants CNS-2006617, CCF-1910958, CCF-2210754, CCF-2210755, CCF-2312508, and CCF-2312509.